# Predicting Polymer Properties Based on Multimodal Multitask Pretraining

### v1

@@ Abstract
In the past few decades, polymers, high-molecular-weight compounds formed by bonding numerous identical or similar monomers covalently, have played an essential role in various scientific fields. In this context, accurate prediction of their properties is becoming increasingly crucial. Typically, the properties of a polymer, such as plasticity, conductivity, bio-compatibility, and so on, are highly correlated with its 3D structure. However, current methods for predicting polymer properties heavily rely on information from polymer SMILES sequences (P-SMILES strings) while ignoring crucial 3D structural information, leading to sub-optimal performance. In this work, we propose MMPolymer, a novel multimodal multitask pretraining framework incorporating both polymer 1D sequential information and 3D structural information to enhance downstream polymer property prediction tasks. Besides, to overcome the limited availability of polymer 3D data, we further propose the "Star Substitution" strategy to extract 3D structural information effectively. During pretraining, MMPolymer not only predicts masked tokens and recovers 3D coordinates but also achieves the cross-modal alignment of latent representation. Subsequently, we further fine-tune the pretrained MMPolymer for downstream polymer property prediction tasks in the supervised learning paradigm. Experimental results demonstrate that MMPolymer achieves state-of-the-art performance in various polymer property prediction tasks. Moreover, leveraging the pretrained MMPolymer and using only one modality (either P-SMILES string or 3D conformation) during fine-tuning can also surpass existing polymer property prediction methods, highlighting the exceptional capability of MMPolymer in polymer feature extraction and utilization. Our online platform for polymer property prediction is available at https://app.bohrium.dp.tech/mmpolymer.
1 Introduction
@@
As illustrated in Figure 1, polymers are constructed by bonding numerous identical or similar monomers based on various polymerization operations (e.g., addition, ring-opening, and condensation) [8]. In the past few decades, polymers have played a crucial role in many scientific fields, such as chemistry [7], material science [6], drug design [33], and bioinformatics [38]. In the applications mentioned above, achieving accurate prediction of polymer properties has garnered more and more attention due to its significance [3, 9]. For example, predicting polymer properties like plasticity and conductivity helps guide the design and development of polymer-based materials with specific functionalities [5]. Besides, predicting polymeric drug carriers’ bio-compatibility and release kinetics is essential for designing effective and safe drugs [34].
Figure 1: An illustration of monomer polymerization and typical polymerization operations.
Traditional studies on polymers rely on experiments or simulations to assess their properties [16, 10]. While these techniques yield precise outcomes, they suffer from poor scalability due to their high costs, extensive time requirements, and the need for specialized tools and knowledge. As a result, they fail to meet the rapidly increasing demands of polymer property prediction. Subsequently, although some learning-based methods [36, 11, 2, 46] have been proposed for polymer property prediction, their performance is often unsatisfactory due to the scarcity of high-quality polymer property data. Recently, some attempts have been made to mitigate the data insufficiency based on model pretraining techniques [53, 20]. However, they merely rely on the information from polymer SMILES sequences (i.e., P-SMILES strings) while ignoring the fact that the properties of a polymer are highly correlated with its 3D structure, which leads to sub-optimal performance as well.
To overcome the limitations of current polymer property prediction methods, we propose a novel multimodal multitask pretraining framework, MMPolymer, leveraging 3D structural information to enhance the predictive capabilities of polymer properties. As illustrated in Figure 2, MMPolymer leverages a two-tower architecture to represent polymer SMILES sequences (i.e., P-SMILES strings) and 3D conformations in the latent space respectively. Aligning latent representations of the two modalities further leads to semantically meaningful polymer representations for downstream polymer property prediction tasks. Notably, considering the regularity of polymer 3D structure and scarcity of polymer 3D data, we leverage the 3D conformation of the corresponding repeating unit  [^1] Not equivalent to the monomer, which lacks polymerization site information. to approximate the whole polymer 3D conformation in our work. Specifically, we first convert the corresponding P-SMILES string through our "Star Substitution" strategy, where the "*" symbol in the P-SMILES string is replaced by the neighboring atom symbol of another "*" symbol. Then we use the RDKit [21] tool to generate the 3D conformation based on the converted P-SMILES string. In this way, the generated 3D conformation can not only reflect 3D structural features within repeating units but also reflect 3D structural features between the repeating units, thus contributing to extracting polymer 3D structural information effectively.
@@
Figure 2: The scheme of the proposed method. Here, the red arrows indicate the pipeline of our multimodal multitask pretraining paradigm, and the blue arrows indicate the pipeline of fine-tuning steps for downstream polymer property prediction tasks. The blue modules are designed for 1D sequences (i.e., P-SMILES strings), and the red modules are designed for 3D conformations. The modules shared by 1D and 3D representations are labeled in green.
@@
During pretraining, we mask the P-SMILES string randomly, add noise to the atom coordinates of corresponding 3D conformation, and pass them through the 1D and 3D representation networks, respectively. We learn the proposed representation networks based on three pretraining tasks, including predicting masked tokens, recovering clear 3D coordinates (i.e., coordinate denoising), and aligning the latent representations across the two modalities by contrastive learning [24]. Solving these three tasks jointly leads to the proposed multimodal multitask pretraining paradigm, which helps our model fully incorporate both polymer 1D sequential and 3D structural information. In the fine-tuning phase, we further fine-tune the pretrained MMPolymer in the supervised learning paradigm for downstream polymer property prediction tasks.
@@
To the best of our knowledge, MMPolymer is the first work that incorporates 3D structural information into polymer property prediction. The experimental results on various polymer property datasets consistently demonstrate that MMPolymer significantly outperforms existing methods, achieving state-of-the-art performance. Moreover, even merely relying on single-modal information (either polymer 1D sequential information or 3D structural information) during fine-tuning, MMPolymer still outperforms existing methods, showcasing its promising capability of extracting and utilizing polymer features. Besides, comprehensive ablation studies offer valuable insights into the performance of our method, further supporting its rationality and effectiveness.
2 Related Work
2.1 Polymer Property Prediction
@@
Polymer property prediction is one of the fundamental problems in polymer informatics, and various machine learning and deep learning algorithms have been widely used as promising tools for polymer property prediction [15], offering efficient alternatives to traditional experimental or simulational approaches [16, 10].
@@
Early works [5, 54] in polymer property prediction employed classic machine learning algorithms (e.g., multiple linear regression) to predict corresponding polymer properties from polymer sequences. The work in [22] further applied kernel support vector machines (kernel SVMs) to deal with non-linear relationships between polymer sequences and their properties, achieving great predictive performance. For deep learning examples, the work in [35] predicted mechanical properties of polymer-carbon nanotube surfaces using convolutional neural networks (CNNs) while recurrent neural networks (RNNs) were also used in [40, 51] to learn the latent relationships between polymer sequences and their properties. However, all these prediction methods are significantly limited by the scarcity of high-quality polymer property data.
Recently, inspired by the exceptional performance of various pretrained models in natural language processing (NLP) tasks [28, 23, 58], some Transformer-based pretraining methods have been proposed for polymer property prediction. These methods like Transpolymer [53] and polyBERT [20] treat polymers as character sequences (i.e., P-SMILES strings) and undergo pretraining on extensive unlabeled polymer sequences, so they heavily rely on the information learned from polymer sequences. However, the latest studies  [41, 59, 49] have revealed that properties are mainly determined by the 3D structure, thus highlighting the crucial need for a paradigm shift towards integrating 3D structural information into property prediction.
2.2 Pretraining for Molecular Modeling
@@
A closely related domain to polymer science is molecular science [26, 13, 42], where numerous pretraining methods have been developed for molecular modeling. These methods have greatly contributed to improving the performance of various downstream tasks in molecular science, especially molecular property prediction.
@@
SMILES-BERT [48] and ChemBERTa [1], were first proposed and achieved great predictive performance on various molecular property prediction tasks by undergoing pretraining on extensive molecular sequences (e.g., SMILES strings). Then, due to the rapid development of graph neural networks [37, 55, 52], subsequent works were further extended to molecular graphs. For example, MolCLR [50] was pretrained on 10 million molecular graphs by contrastive learning strategy, and the work in [4] pretrained graph neural networks by utilizing inherent properties of molecular graphs.
Recently, considering that intramolecular interactions are fundamentally three-dimensional [12, 45, 61], many works tend to incorporate molecular 3D structural information into pretraining to acquire more comprehensive molecular representation. For example, the work in [60] proposed a unified 2D and 3D pretraining framework for more informative representation, while 3D Infomax [41] proposed to maximize the mutual information between learned 3D representation and 2D representation during pretraining to improve the performance of molecular property prediction. Furthermore, several 3D pretraining methods [59, 49] have been proposed and demonstrated significant performance gains across various downstream tasks in molecular science, thus underscoring the critical role of 3D structural information.
3 Method
3.1 Overview
@@
To achieve our goal that acquiring a comprehensive polymer representation for downstream polymer property prediction tasks, we pretrain MMPolymer through the multimodal multitask paradigm. As shown in Figure 2, the 1D representation network $f^{1d}:\mathcal{S}\mapsto\mathcal{X}$ is pretrained via the masked prediction task. Meanwhile, to effectively incorporate 3D structural information, the 3D representation network $f^{3d}:\mathcal{C}\mapsto\mathcal{X}$ is pretrained via the coordinate denoising task. Besides, we further align the latent representations across the two modalities via the cross-modal alignment task during pretraining. Here, $\mathcal{S}$ is the collection of P-SMILES strings, $\mathcal{C}$ is the collection of 3D conformations, and $\mathcal{X}\subset\mathbb{R}^{d}$ is the latent space.
3.2 Model Architecture
3.2.1 1D Representation Network
The 1D representation network $f^{1d}$, based on Transformer architecture [44, 25], takes polymer sequences as input, and outputs the corresponding 1D sequential representation $\bm{X}_{1d}$, aiming to encode as much information as possible about polymer sequences.
Unlike small molecules, which can be easily represented by corresponding sequences (e.g., SMILES string), converting polymers to sequences is not straightforward due to their complex structures and compositions. Here, we use the P-SMILES string [17], a modified SMILES string for polymers, as input polymer sequences. Specifically, the P-SMILES string is formed by combining the SMILES string [32] of the corresponding monomer with two "*" symbols which are used to indicate the polymerization sites (i.e., the connected atoms between repeating units). For example, polyethylene is represented by "*CC*" and polypropylene is represented by "*CC(*)C". Besides, if other descriptors like the polymerization degree are also available, they can be added to the end of the corresponding P-SMILES string and fed into our 1D representation network together.
Figure 3: Left: the architecture of our 1D representation network, which takes polymer SMILES sequences (i.e., P-SMILES strings) as input, and outputs corresponding 1D sequential representation. Right: the architecture of our 3D representation network, which takes 3D conformation as input, and outputs corresponding 3D structural representation.
As shown in Figure 3, given a polymer sequence represented by the corresponding P-SMILES string "*CC(*)F", we first use chemical-aware tokenization from  [53] to get corresponding tokens. Here, the P-SMILES string "*CC(*)F" is converted into [‘*’, ‘C’, ‘C’, ‘(’, ‘*’, ‘)’, ‘F’] along with special tokens (i.e., [CLS] and [SEP]). After tokenization, these tokens are further converted into continuous vector representations by adding each token’s embedding to its corresponding position embedding. As a result, these vectorized tokens can be effectively interpreted and handled by the attention-based backbone of our 1D representation network. Specifically, segment embeddings are removed and we use the typical trigonometric function to acquire corresponding position embedding, i.e.,
@@ 3.2 Model Architecture
Finally, the embedding vectors are fed into the attention-based backbone of our 1D representation network to get the corresponding 1D sequential representation $\bm{X}_{1d}$. Here we utilize the output [CLS] representation as the corresponding 1D sequential representation $\bm{X}_{1d}$ for the input polymer sequence, i.e.,
$\bm{X}_{1d}=Rep_{[CLS]}.$ (3)
3.2.2 3D Representation Network
The 3D representation network $f^{3d}$, based on SE(3)-Transformer architecture [14, 59], takes 3D conformation as input and outputs corresponding SE(3)-invariant representation vector as 3D structural representation $\bm{X}_{3d}$, aiming to encode comprehensive polymer 3D structural information.
Figure 4: Visualization of our "Star Substitution" strategy.
"Star Substitution" strategy. Notably, due to the regularity of polymer 3D structure and scarcity of polymer 3D data, we approximate the whole polymer 3D conformation by using the 3D conformation of its corresponding repeating unit. Specifically, as shown in Figure 4, we first generated the converted P-SMILES string through our "Star Substitution" strategy, where the "*" symbol in the P-SMILES string is replaced by the neighboring atom symbol of another "*" symbol. For example, "*NCCCCCC(*)=O" is converted to "CNCCCCCC(N)=O" and "*Oc1ccc(CC(*)=O)cc1" is converted to "COc1ccc(CC(O)=O)cc1". Then we further generate the corresponding 3D conformation by RDKit [21] based on the converted P-SMILES string. In this way, the generated 3D conformation can not only reflect 3D structural features within repeating units but also reflect 3D structural features between the repeating units, thus contributing to approximating the whole polymer 3D conformation as much as possible.
@@
Meanwhile, a pair-type aware Gaussian kernel [39] encodes the Euclidean distances of each atom pair $(a_{i},a_{j})$ as the initial pair-level representation $\bm{X}^{(0)}_{p}\in\mathbb{R}^{N\times N\times d_{p}}$, i.e.,
$\bm{X}^{(0)}_{p}[i,j]=\mathcal{G}(\bm{v}_{{a_{i}}{a_{j}}}\|\bm{p}_{i}-\bm{p}_{j}\|+\bm{u}_{{a_{i}}{a_{j}}}-\bm{\mu},\bm{\sigma}).$ (5)
@@
Then, the initial representation $\bm{X}^{(0)}_{a}$ and $\bm{X}^{(0)}_{p}$ are fed into the attention-based backbone of our 3D representation network to get the final representation $\bm{X}^{(L)}_{a}$ and $\bm{X}^{(L)}_{p}$ after passing $L$ encoder layers. To fully capture intricate 3D structure information, the multi-head self-attention mechanism is applied to the encoder layers, and the number of attention heads is determined by the dimension of pair-level embedding $d_{p}$. Besides, to update atom-level and pair-level representation effectively, we use atom-to-pair communication to further improve the traditional self-attention mechanism. The specific implementation of the $h$-th attention head in the $l$-th encoder layer can be expressed as follows:
$\bm{Atten}^{(l,h)}=\frac{\bm{Q}^{(l,h)}(\bm{K}^{(l,h)})^{\top}}{\sqrt{d_{h}}}+\bm{X}^{(l-1,h)}_{p},$ (6)
$\bm{X}^{(l,h)}_{a}=\sigma(\bm{Atten}^{(l,h)})\bm{V}^{(l,h)},\quad\bm{X}^{(l,h)}_{p}=\bm{Atten}^{(l,h)},$ (7)
@@
$\bm{X}^{(l)}_{p}=\text{Concat}(\{\bm{X}^{(l,h)}_{p}\}_{h=1}^{d_{p}}),$ (8)
$\bm{X}^{(l)}_{a}=\text{Concat}(\{\bm{X}^{(l,h)}_{a}\}_{h=1}^{d_{p}})\bm{W}_{O}^{(l)}+\bm{X}^{(l-1)}_{a},$ (9)
@@
$\bm{X}_{3d}=\bm{X}^{(L)}_{a}[0,:].$ (10)
3.3 Learning Paradigm
As shown in Figure 2, we pretrain our MMPolymer with a multimodal multitask paradigm, including the masked prediction task for pretraining the 1D representation network $f^{1d}$ based on polymer 1D sequence data in Sec. 3.3.1, the coordinate denoising task for pretraining the 3D representation network $f^{3d}$ based on polymer 3D structure data in Sec. 3.3.2, and the cross-modal alignment task for aligning the learned latent representations across the two modalities in Sec. 3.3.3, to acquire a comprehensive polymer representation for downstream polymer property prediction tasks. The total loss during pretraining is expressed as follows:
$\mathcal{L}_{\text{pretrain}}=\mathcal{L}_{\text{1D}}+\mathcal{L}_{\text{3D}}+\mathcal{L}_{\text{contrastive}}.$ (11)

### v2

# MMPolymer: A Multimodal Multitask Pretraining Framework for Polymer Property Prediction

@@ Abstract.
Polymers are high-molecular-weight compounds constructed by the covalent bonding of numerous identical or similar monomers so that their 3D structures are complex yet exhibit unignorable regularity. Typically, the properties of a polymer, such as plasticity, conductivity, bio-compatibility, and so on, are highly correlated with its 3D structure. However, existing polymer property prediction methods heavily rely on the information learned from polymer SMILES sequences (P-SMILES strings) while ignoring crucial 3D structural information, resulting in sub-optimal performance. In this work, we propose MMPolymer, a novel multimodal multitask pretraining framework incorporating polymer 1D sequential and 3D structural information to encourage downstream polymer property prediction tasks. Besides, considering the scarcity of polymer 3D data, we further introduce the "Star Substitution" strategy to extract 3D structural information effectively. During pretraining, in addition to predicting masked tokens and recovering clear 3D coordinates, MMPolymer achieves the cross-modal alignment of latent representations. Then we further fine-tune the pretrained MMPolymer for downstream polymer property prediction tasks in the supervised learning paradigm. Experiments show that MMPolymer achieves state-of-the-art performance in downstream property prediction tasks. Moreover, given the pretrained MMPolymer, utilizing merely a single modality in the fine-tuning phase can also outperform existing methods, showcasing the exceptional capability of MMPolymer in polymer feature extraction and utilization.
Multimodal Multitask Pretraining, Polymer Property Prediction, Polymer Informatics, Bioinformatics
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USAbooktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAdoi: 10.1145/3627673.3679684isbn: 979-8-4007-0436-9/24/10ccs: Computing methodologies Multi-task learningccs: Information systems Data mining
1. Introduction
@@
In the past few decades, polymers have played a crucial role in many scientific fields, such as chemistry (Carraher Jr, 2017), material science (Brazel and Rosen, 2012), drug design (Pasut and Veronese, 2007), and bioinformatics (Schaffert and Wagner, 2008). In the applications mentioned above, achieving accurate prediction of polymer properties has garnered more and more attention due to its significance (Audus and de Pablo, 2017; Chen et al., 2021). For example, predicting polymer properties like plasticity and conductivity helps guide the design and development of polymer-based materials with specific functionalities (Bicerano, 2002). Besides, predicting polymeric drug carriers’ bio-compatibility and release kinetics is essential for designing effective and safe drugs (Qiu and Bae, 2006).
@@
Figure 1. The scheme of the proposed method. Here, the red arrows indicate the pipeline of our multimodal multitask pretraining paradigm, and the blue arrows indicate the pipeline of fine-tuning steps for downstream polymer property prediction tasks. The blue modules are designed for 1D sequences (i.e., P-SMILES strings), and the red modules are designed for 3D conformations. The modules shared by 1D and 3D representations are labeled in green.
Traditional studies on polymers rely on experiments or simulations to assess their properties (Huan and Ramprasad, 2020; Danielsen et al., 2021). While these techniques yield precise outcomes, they suffer from poor scalability due to their high costs, extensive time requirements, and the need for specialized tools and knowledge. As a result, they fail to meet the rapidly increasing demands of polymer property prediction. Subsequently, although some learning-based methods (Roy et al., 2006; Doan Tran et al., 2020; Amamoto, 2022; Wang et al., 2023a) have been proposed for polymer property prediction, their performance is often unsatisfactory due to the scarcity of high-quality polymer property data. Recently, some attempts have been made to mitigate data insufficiency based on pretraining techniques (Xu et al., 2023; Kuenneth and Ramprasad, 2023). However, these attempts merely rely on the information from polymer SMILES sequences (i.e., P-SMILES strings) while ignoring the fact that the properties of a polymer are highly correlated with its 3D structure, which leads to sub-optimal performance as well.
To overcome the limitations of existing polymer property prediction methods, we propose a novel multimodal multitask pretraining framework, MMPolymer, leveraging 3D structural information to enhance the predictive capabilities of polymer properties. Here, considering the regularity of polymer 3D structure and scarcity of polymer 3D data, we leverage the 3D conformation of the corresponding repeating unit [^1] Not equivalent to the monomer, which lacks polymerization site information. to approximate the whole polymer 3D conformation. Specifically, we first convert the corresponding P-SMILES string through our "Star Substitution" strategy, where the "*" symbol in the P-SMILES string is replaced by the neighboring atom symbol of another "*" symbol. Then we use the RDKit (Landrum et al., 2013) tool to generate the 3D conformation based on the converted P-SMILES string. In this way, the generated 3D conformation can not only reflect 3D structural features within repeating units but also reflect 3D structural features between the repeating units, thus contributing to extracting polymer 3D structural information effectively.
@@
During pretraining, we mask the P-SMILES string randomly, add noise to the atom coordinates of corresponding 3D conformation, and pass them through the 1D and 3D representation networks, respectively. As illustrated in Figure 1, we learn the proposed representation networks based on three pretraining tasks, including predicting masked tokens, recovering clear 3D coordinates (i.e., coordinate denoising), and aligning the latent representations across the two modalities by contrastive learning (Liu et al., 2021). Solving these three tasks jointly leads to the proposed multimodal multitask pretraining paradigm, which helps our model fully incorporate polymer 1D sequential and 3D structural information. Then we further fine-tune the pretrained MMPolymer in the supervised learning paradigm for downstream polymer property prediction tasks.
@@
To the best of our knowledge, MMPolymer is the first work that incorporates 3D structural information into polymer property prediction. The experimental results on various polymer property datasets consistently demonstrate that MMPolymer significantly outperforms existing polymer property prediction methods, achieving state-of-the-art performance. Moreover, even merely relying on single-modal information (either polymer 1D sequential or 3D structural information) during fine-tuning, MMPolymer still outperforms existing polymer property prediction methods, showcasing its promising capability of extracting and utilizing polymer features. Besides, comprehensive ablation studies are also conducted to offer valuable insights into the performance of MMPolymer, further supporting its rationality and effectiveness.
Figure 2. Left: the architecture of our 1D representation network, which takes polymer SMILES sequences (i.e., P-SMILES strings) as input, and outputs corresponding 1D sequential representation. Right: the architecture of our 3D representation network, which takes 3D conformation as input, and outputs corresponding 3D structural representation.
2. Related Work
2.1. Polymer Property Prediction
@@
Polymer property prediction is one of the fundamental problems in polymer informatics, and various machine learning and deep learning algorithms have been widely used as promising tools (Hatakeyama-Sato, 2023).
@@
Early works (Bicerano, 2002; Yu et al., 2006) in polymer property prediction employed classic machine learning algorithms (e.g., multiple linear regression) to predict corresponding polymer properties from polymer sequences. The work in (Le et al., 2012) further applied kernel support vector machines (kernel SVMs) to deal with non-linear relationships between polymer sequences and their properties, achieving great predictive performance. For deep learning examples, the work in (Rahman et al., 2021) predicted mechanical properties of polymer-carbon nanotube surfaces using convolutional neural networks (CNNs) while recurrent neural networks (RNNs) were also used in (Simine et al., 2020; Webb et al., 2020) to learn the latent relationships between polymer sequences and their properties. However, these methods are significantly limited by the scarcity of high-quality polymer property data. Recently, inspired by the exceptional performance of various pretrained models in natural language processing (NLP) tasks (Min et al., 2023; Liu et al., 2023; Zhao et al., 2023), some Transformer-based pretraining methods have been proposed for polymer property prediction. These methods like Transpolymer (Xu et al., 2023) and polyBERT (Kuenneth and Ramprasad, 2023) treat polymers as character sequences (i.e., P-SMILES strings) and undergo pretraining on extensive unlabeled polymer sequences, so they heavily rely on the information learned from polymer sequences. However, the latest studies  (Stärk et al., 2022; Zhou et al., 2023; Wang et al., 2023c) have revealed that properties are mainly determined by the 3D structure, thus highlighting the crucial need for a paradigm shift towards integrating 3D structural information into polymer property prediction.
2.2. Pretraining for Molecular Modeling
@@
A closely related domain to polymer science is molecular science (Ma et al., 2022; Feng et al., 2022; Sun et al., 2022), where numerous pretraining methods have been developed for molecular modeling in the past few years.
@@
SMILES-BERT (Wang et al., 2019) and ChemBERTa (Ahmad et al., 2022) were first proposed and achieved great predictive performance on various molecular property prediction tasks by undergoing pretraining on extensive molecular sequences (e.g., SMILES strings). Then, due to the rapid development of graph neural networks (Satorras et al., 2021; Yuan et al., 2021; Wu et al., 2022), subsequent works were extended to molecular graphs. For example, MolCLR (Wang et al., 2022) was pretrained on 10 million molecular graphs by contrastive learning strategy, and the work in (Benjamin et al., 2022) pretrained graph neural networks by utilizing inherent properties of molecular graphs. Recently, considering that intramolecular interactions are fundamentally three-dimensional (Fang et al., 2022; Wang et al., 2023b; Zhu et al., 2023), many works tend to incorporate 3D structural information into pretraining to acquire more comprehensive molecular representation. For example, the work in (Zhu et al., 2022) proposed a unified 2D and 3D pretraining framework for more informative representation, and 3D Infomax (Stärk et al., 2022) proposed to maximize the mutual information between 3D and 2D representation during pretraining to improve the performance of molecular property prediction.
3. Method
3.1. Overview
@@
To acquire a comprehensive polymer representation for downstream polymer property prediction tasks, we pretrain MMPolymer through the multimodal multitask paradigm. As shown in Figure 1, the 1D representation network $f^{1d}:\mathcal{S}\mapsto\mathcal{X}$ is pretrained via the masked prediction task. Meanwhile, to effectively incorporate 3D structural information, the 3D representation network $f^{3d}:\mathcal{C}\mapsto\mathcal{X}$ is pretrained via the coordinate denoising task. Besides, we further align the latent representations across the two modalities via the cross-modal alignment task during pretraining. Here, $\mathcal{S}$ is the collection of P-SMILES strings, $\mathcal{C}$ is the collection of 3D conformations, and $\mathcal{X}\subset\mathbb{R}^{d}$ is the latent space.
3.2. Model Architecture
3.2.1. 1D Representation Network
The 1D representation network $f^{1d}$, based on Transformer architecture (Vaswani et al., 2017), takes polymer sequences as input, and outputs the corresponding 1D sequential representation $\bm{X}_{1d}$, aiming to encode as much information as possible about polymer sequences.
Unlike small molecules, which can be easily represented by corresponding sequences (e.g., SMILES string), converting polymers to sequences is not straightforward due to their complex structures and compositions. Here, we use the P-SMILES string (Kim et al., 2018), a modified SMILES string for polymers, as input polymer sequences. Specifically, the P-SMILES string is formed by combining the SMILES string (O’Boyle, 2012) of the corresponding monomer with two "*" symbols which are used to indicate the polymerization sites (i.e., the connected atoms between repeating units). For example, polyethylene is represented by "*CC*" and polypropylene is represented by "*CC(*)C". Besides, if other descriptors (e.g., polymerization degree) are also available, they can be added to the end of the corresponding P-SMILES string and fed into our 1D representation network together.
Figure 3. Visualization of our "Star Substitution" strategy.
As shown in Figure 2, given a polymer sequence represented by the corresponding P-SMILES string "*CC(*)F", we first use chemical-aware tokenization from  (Xu et al., 2023) to get corresponding tokens. Here, the P-SMILES string "*CC(*)F" is converted into [‘*’, ‘C’, ‘C’, ‘(’, ‘*’, ‘)’, ‘F’] along with special tokens (i.e., [CLS] and [SEP]). After tokenization, these tokens are further converted into continuous vector representations by adding each token’s embedding to its corresponding position embedding. As a result, these vectorized tokens can be effectively interpreted and handled by the attention-based backbone of our 1D representation network. Specifically, segment embeddings are removed and we use the typical trigonometric function to acquire corresponding position embedding, i.e.,
@@ 3.2. Model Architecture
Finally, the embedding vectors are fed into the attention-based backbone (the same as the one used in Roberta (Liu et al., 2019)) of our 1D representation network to get the corresponding 1D sequential representation $\bm{X}_{1d}$. Here we utilize the output [CLS] representation as the corresponding 1D sequential representation $\bm{X}_{1d}$ for the input polymer sequence, i.e.,
(3) $\bm{X}_{1d}=Rep_{[CLS]}.$
3.2.2. 3D Representation Network
The 3D representation network $f^{3d}$, based on SE(3)-Transformer architecture (Fuchs et al., 2020), takes 3D conformation as input and outputs corresponding SE(3)-invariant representation vector as 3D structural representation $\bm{X}_{3d}$, aiming to encode comprehensive polymer 3D structural information.
"Star Substitution" strategy. Due to the regularity of polymer 3D structure and scarcity of polymer 3D data, we approximate the whole polymer 3D conformation with the 3D conformation of its corresponding repeating unit. Specifically, as shown in Figure 3, we first generated the converted P-SMILES string through our "Star Substitution" strategy, where the "*" symbol in the P-SMILES string is replaced by the neighboring atom symbol of another "*" symbol. For example, "*NCCCCCC(*)=O" is converted to "CNCCCCCC(N)=O" and "*Oc1ccc(CC(*)=O)cc1" is converted to "COc1ccc(CC(O)=O)cc1". Then we further generate the corresponding 3D conformation by RDKit (Landrum et al., 2013) based on the converted P-SMILES string. In this way, the generated 3D conformation can not only reflect 3D structural features within repeating units but also reflect 3D structural features between the repeating units, thus contributing to approximating the whole polymer 3D conformation as much as possible.
@@
Meanwhile, a pair-type aware Gaussian kernel (Scholkopf et al., 1997) encodes the Euclidean distances of each atom pair $(a_{i},a_{j})$ as the initial pair-level representation $\bm{X}^{(0)}_{p}\in\mathbb{R}^{N\times N\times d_{p}}$, i.e.,
(5) $\bm{X}^{(0)}_{p}[i,j]=\mathcal{G}(\bm{v}_{{a_{i}}{a_{j}}}\|\bm{p}_{i}-\bm{p}_{j}\|+\bm{u}_{{a_{i}}{a_{j}}}-\bm{\mu},\bm{\sigma}).$
@@
Then, the initial representation $\bm{X}^{(0)}_{a}$ and $\bm{X}^{(0)}_{p}$ are fed into the attention-based backbone (the same as the one used in Uni-Mol (Zhou et al., 2023)) of our 3D representation network to get the final representation $\bm{X}^{(L)}_{a}$ and $\bm{X}^{(L)}_{p}$. To fully capture intricate 3D structure information, the multi-head self-attention mechanism is applied to the encoder layers, and the number of attention heads is determined by the dimension of pair-level embedding $d_{p}$. Besides, to update atom-level and pair-level representation effectively, we use atom-to-pair communication to improve the traditional self-attention mechanism. The specific implementation of the $h$-th attention head in the $l$-th encoder layer can be expressed as follows:
(6) $\bm{Atten}^{(l,h)}=\frac{\bm{Q}^{(l,h)}(\bm{K}^{(l,h)})^{\top}}{\sqrt{d_{h}}}+\bm{X}^{(l-1,h)}_{p},$
(7) $\bm{X}^{(l,h)}_{a}=\sigma(\bm{Atten}^{(l,h)})\bm{V}^{(l,h)},\quad\bm{X}^{(l,h)}_{p}=\bm{Atten}^{(l,h)},$
@@
(8) $\bm{X}^{(l)}_{a}=\text{Concat}(\{\bm{X}^{(l,h)}_{a}\}_{h=1}^{d_{p}})\bm{W}_{O}^{(l)}+\bm{X}^{(l-1)}_{a},$
(9) $\bm{X}^{(l)}_{p}=\text{Concat}(\{\bm{X}^{(l,h)}_{p}\}_{h=1}^{d_{p}}),$
@@
(10) $\bm{X}_{3d}=\bm{X}^{(L)}_{a}[0,:].$
3.3. Learning Paradigm
As shown in Figure 1, we pretrain our MMPolymer with a multimodal multitask paradigm, including the masked prediction task for pretraining the 1D representation network $f^{1d}$ based on polymer 1D sequence data in Sec. 3.3.1, the coordinate denoising task for pretraining the 3D representation network $f^{3d}$ based on polymer 3D structure data in Sec. 3.3.2, and the cross-modal alignment task for aligning the learned latent representations across the two modalities in Sec. 3.3.3, to acquire a comprehensive polymer representation for downstream polymer property prediction tasks. The total loss during pretraining can be expressed as follows:
(11) $\mathcal{L}_{\text{pretrain}}=\mathcal{L}_{\text{1D}}+\mathcal{L}_{\text{3D}}+\mathcal{L}_{\text{contrastive}}.$